# The TRGB--SBF Project. III. Refining the HST Surface Brightness Fluctuation Distance Scale Calibration with JWST

### v1

@@ Abstract
The TRGB--SBF Project team is developing an independent distance ladder using a geometrical calibration of the tip of the red giant branch (TRGB) method in elliptical galaxies that can in turn be used to set the surface brightness fluctuation (SBF) distance scale independent of Cepheid variables and Type Ia supernovae. In this paper, we use JWST TRGB distances calibrated using the megamaser galaxy NGC 4258 to determine a new Cepheid-independent SBF zero point for HST. This new calibration, along with improved optical color measurements from PanSTARRS and DECam, gives an updated value of the Hubble-Lemaître constant $H_{0}=73.8\pm 0.7\pm 2.3$ km s-1 Mpc-1 that is virtually identical to the previously published result based on SBF distances calibrated using Cepheids (Blakeslee et al. 2021). Future improvements in the geometrical calibration of TRGB and additions to the HST SBF survey will further reduce the random and systematic uncertainties on $H_{0}$ measured using the extensive HST WFC3/IR SBF dataset. JWST observations of the Coma cluster will soon establish a foundation for the SBF calibration that will extend to much larger distances than is possible with HST.
@@ 1 Introduction
The current disagreement between the measured expansion rate of the universe and that computed from the standard $\Lambda$CDM cosmological model tied to the cosmic microwave background fluctuations is now greater than 5-$\sigma$ (Planck Collaboration et al., 2020; Riess et al., 2022). Resolving this discrepancy is the most pressing issue in cosmology today. It is not yet clear if the explanation is an accumulation of various systematic errors in local distance measurements or if new physics beyond the $\Lambda$CDM model is required (e.g., Di Valentino et al., 2021).
@@
The TRGB--SBF Project seeks to reduce or eliminate systematic uncertainties in the first rungs of the distance ladder by developing a new distance scale based on a geometrical calibration of old, metal-rich populations in massive elliptical galaxies (i.e., Pop II RGB stars) that is completely independent of the Cepheid and Type Ia Supernova distance measurements. To do this, we are connecting Tip of the Red Giant Branch (TRGB) and Surface Brightness Fluctuation (SBF) distances measured in 14 nearby elliptical galaxies using the James Webb Space Telescope (Tully et al., 2023). The TRGB--SBF Project will eventually link geometrically-calibrated TRGB distances at 15-20 Mpc with SBF distances reaching out to perhaps 300 Mpc or more using the same telescope (JWST), camera (NIRCam), and filters to minimize systematic uncertainties for each rung of the ladder.
@@
SBF distances have already been measured in more than 370 galaxies out to about 100 Mpc using the Hubble Space Telescope (HST), including more than 220 in F110W using the WFC3/IR camera. In this paper, we use the TRGB distances to eight individual galaxies calibrated using the geometrical distance to the megamaser galaxy NGC 4258 (Reid et al., 2019; Anand et al., 2024a) to calibrate the HST SBF distances via TRGB. In the future, the Gaia satellite (Gaia Collaboration et al., 2017, 2018) will add additional geometrical anchors to the TRGB distance scale to further reduce the systematic calibration uncertainty. This study presents the first step towards an SBF calibration that is independent of the Cepheid distance scale and connects JWST TRGB to the existing HST SBF distance measurements out to 100 Mpc.
@@ 2 Infrared Surface Brightness Fluctuations with HST
The current generation of optical and IR HST SBF measurements are based on the calibrations published by Mei et al. (2007); Blakeslee et al. (2009, 2010) and Jensen et al. (2015, 2021). The foundation of all the HST SBF measurements to-date is the HST Key Project Cepheid distances to seven galaxies with ground-based SBF distances measured by Tonry et al. (2001) and Blakeslee et al. (2002), combined with a large number of overlapping ground-based and HST SBF measurements in the Virgo and Fornax galaxy clusters (see Blakeslee et al., 2010, for direct comparisons). These large surveys provided the statistical foundation for SBF and showed that the intrinsic scatter in the SBF technique is  ${\sim\,}0.06$ mag (optical) to  ${\sim\,}0.1$ mag (near-IR) after correction for stellar population variations using optical colors (Blakeslee et al., 2009; Jensen et al., 2015).
@@
Fluctuations are much brighter at near-IR wavelengths than in the optical and are measurable to ${\gtrsim}80$ Mpc in a single HST orbit. SBF is much more efficient than Cepheids or supernovae for distances out to 100 Mpc and achieves comparable precision. The majority of the SBF distances beyond 40 Mpc useful for measuring $H_{0}$ were made using NICMOS (Jensen et al., 2001) and WFC3/IR (Cantiello et al., 2018; Blakeslee et al., 2021). The most recent WFC3/IR F110W SBF calibration was determined using 11 galaxies in the Fornax and Virgo galaxy clusters, and was corrected for stellar population variations using ACS $(g_{475}{-}z_{850})$ colors (Jensen et al., 2015). Updates to the Jensen et al. (2015) calibration to include better Galactic foreground extinction corrections and an improved distance to the LMC were made as described by Cantiello et al. (2018) and Jensen et al. (2021).
@@
The majority of the HST IR SBF observations published between 2015 and 2021 were part of HST programs GO-14219 (PI J. Blakeslee), GO-14654 (PI P. Milne), and GO-15625 (PI J. Blakeslee). Jensen et al. (2021) published SBF distances to 63 galaxies from these programs, including 23 supernova host galaxies. Three more HST programs currently nearing completion, GO-16262 (PI R. Tully), GO-17436 (PI J. Jensen), and GO-17446 (PI P. Milne) have already observed ${\gtrsim}164$ more galaxies out to ${\gtrsim}85$ Mpc, a sample that more than doubles the number of type Ia supernova host galaxies with high-quality SBF distances (now 51). To make the most of these new datasets, in this paper we update the HST SBF calibration using the new JWST TRGB distances.
@@
The SBF distances published by Jensen et al. (2021) were used by Blakeslee et al. (2021) to measure a value of the Hubble-Lemaître constant of $H_{0}=73.3\pm 0.7$ (statistical) ±$\pm 2.4$ (systematic)  km s-1 Mpc-1, using both Cepheid and TRGB calibrations to help reduce systematic uncertainties. The subsample of supernova host galaxies provided an alternate connection between Cepheids and Ia SN for massive early-type galaxies via SBF: Garnavich et al. (2023) measured a value of $H_{0}=74.6\pm 0.9\pm 2.7$  km s-1 Mpc-1 using SBF distances and SN light curve measurements from Scolnic et al. (2022). When the SALT2 fit parameters were re-derived by Garnavich et al. (2023) for the fast-declining SNe typically found in the early-type galaxies, the value of $H_{0}$ became $73.3\pm 1.0\pm 2.7$ km s-1 Mpc-1.
@@ 3 Re-calibrating the HST WFC3/IR SBF Distance Scale using TRGB
JWST observations of the TRGB in 14 nearby elliptical galaxies (GO-3055; Tully et al. 2023) provide a new and direct calibration of IR SBF based on the geometrical megamaser distance to NGC 4258 (Anand et al., 2024a; Reid et al., 2019). We will eventually use JWST to determine $H_{0}$ directly by measuring SBF distances out to  ${\sim}\,300$ Mpc, but building a database of distant SBF measurements will take time, and may never reach the quantities already observed with HST over the last 30 years (more than $370$ galaxies, including SBF distances from NICMOS, ACS, and WFC3/IR). To leverage the full potential of the HST database, we need to connect the new JWST TRGB distances to the existing HST SBF dataset. Three of the JWST TRGB target galaxies in the Fornax cluster and five in Virgo have also been observed with WFC3/IR, as shown in Table 1. The new JWST TRGB measurements of these galaxies (Anand et al., 2024b, 2025) provide us with an opportunity to reduce the systematic uncertainties in the SBF distances originating in the Cepheid calibration. Using a direct geometrical calibration of TRGB from NGC 4258 (Reid et al., 2019; Anand et al., 2024a) and SBF measurements in the same galaxies, using the same telescope, camera, and filters, we can eliminate some sources of uncertainty (e.g., Cepheid metallicity corrections in different galaxy types and positional offsets between spirals and ellipticals in the calibration clusters).
@@ 3.1 Updated Calibration
The most direct re-calibration of the HST IR SBF distance scale is to replace the distances from Jensen et al. (2015) with the new JWST-based TRGB distances to the eight galaxies in common between the HST and JWST SBF calibration data sets. The SBF calibration of absolute fluctuation magnitude $\overline{M}_{110}$ depends on the distance scale zero point and a slope with galaxy color, which corrects $\overline{M}_{110}$ for variations in stellar population age and metallicity (e.g., Jensen et al., 2003). Jensen et al. (2015) found that the calibration scatter was lowest when using the average distance to the cluster instead of individual optical SBF distances. We can do the same here, and by using the mean cluster distances from the JWST TRGB measurements we include four more galaxies in Fornax and two more in Virgo, for a total of 14 galaxies in our new calibration (see Table 1).
@@
footnotetext: (4) SBF magnitudes and optical colors have been corrected for Galactic extinction using Schlafly & Finkbeiner (2011) and the NED extinction calculator https://doi.org/10.26132/NED5 (catalog DOI: 10.26132/NED5). All fluctuation magnitudes and colors were measured in two annular region around each galaxy from 8.2 to 16.4 arcsec and 16.4 to 32.8 arcsec in radius. The values shown are the weighted averages of the two measurements.
@@
The direct overlap between JWST and HST calibrations includes three galaxies in the Fornax cluster and five in Virgo (including NGC 4636 and NGC 4697 in peripheral in-falling regions). Using the combined average distance to Virgo from TRGB (Anand et al., 2024b, 2025) and the ACSVCS survey average distance modulus of $31.092\pm 0.013$ mag (Mei et al., 2007; Blakeslee et al., 2009), which was used to calibrate the previous IR SBF measurements, the weighted mean difference is $\Delta\mu_{TRGB-SBF}=-0.067\pm 0.031$ mag (Anand et al., 2025). For Fornax, the difference between the mean ACS SBF distance modulus of 31.51 mag (Blakeslee et al., 2009) used for the Jensen et al. (2015) IR SBF calibration and the mean distance from three JWST TRGB measurements gives a difference of $\Delta\mu=-0.086\pm 0.077$ mag, where the error results from the scatter among the three galaxies.
@@
The SBF measurements were updated since the original calibration published by Jensen et al. (2015), and for this study we have included three additional calibration galaxies (NGC 4552, NGC 4636, and NGC 4649). The new $\overline{m}_{110}$ measurements for the calibrators are consistent with the procedures we used to measure $\overline{m}_{110}$ in the sample from which we derived $H_{0}$ (Jensen et al., 2021; Blakeslee et al., 2021). The average difference between the SBF magnitudes $\overline{m}_{110}$ from the Jensen et al. (2015) calibration paper and the present values for the five calibrator galaxies in common is 0.072 mag with a standard deviation of 0.03 mag. There have been a number of procedural improvements to the SBF analysis that were previously applied to the sample used to determine $H_{0}$ (Blakeslee et al., 2021), and those have been applied to these new measurements. We previously identified an offset of $+0.05\pm 0.02$ mag to the 2015 calibration values due to the PSF normalization (Cantiello et al., 2018). Given the small sample size in this study, the 0.072 mag offset is considered to be statistically consistent with the previously measured 0.05 mag offset. Further assessment of the reliability of the $\overline{m}_{110}$ measurements will require additional overlap between calibration samples.
@@
Optical galaxy colors are critical for the SBF method as they are necessary to correct the SBF amplitudes for intrinsic variations in the ages and metallicities of the stellar populations in elliptical galaxies (Jensen et al., 2003; Blakeslee et al., 2009; Jensen et al., 2015). The original IR SBF calibration was based on the ACS $(g_{475}{-}z_{850})$ colors, which are not generally available for the vast majority of the WFC3/IR observations in the HST archive. Jensen et al. (2021) transformed PanSTARRS $(g{-}z)$ colors (Chambers et al., 2019) to the ACS photometric frame for computing distances. For this study, we remeasured all the PanSTARRS $(g{-}z)$ colors and determined the SBF calibration directly in that system. We found that for galaxies with significant color gradients, the PanSTARRS $(g{-}z)$ colors used by Jensen et al. (2021) for computing SBF distances—measured in the high-$S/N$ ratio central regions—did not always match the colors in the regions where the SBF was measured. For typical massive galaxies with slightly redder cores and bluer halos, correcting the colors to match the SBF regions leads to larger distances, by as much as 5% in some cases, and hence this revision tends to reduce the value of $H_{0}$. For the 61 galaxies in the Jensen et al. (2021) dataset with PanSTARRS photometry, which are used again in this paper to recompute $H_{0}$, the color shift was measured to be $-0.023\pm 0.019$ mag (bluer). The average photometric uncertainty in $(g{-}z)$ is 0.02 mag for PanSTARRS, which includes the uncertainty in the extinction correction (Schlafly & Finkbeiner, 2011) added in quadrature. Given that the historical slope of the relationship between SBF $\overline{M}_{110}$ and $(g{-}z)$ is  ${\sim\,}2$ (Jensen et al., 2015), the expected shift in distance modulus is approximately 0.05 mag or 2.3% in distance, counteracting in part the increase in $H_{0}$ expected from the new TRGB zero point.
@@
The original SBF calibration of Jensen et al. (2015) included equal numbers of Fornax and Virgo cluster galaxies, which were calibrated using ACS optical photometry. The JWST TRGB sample includes three galaxies in the Fornax cluster. The PanSTARRS-1 survey does not reach far enough south to include Fornax, so the current TRGB calibration is limited to five galaxies in Virgo. DECam survey provides a higher-$S/N$ alternative to the PanSTARRS calibration in both Fornax and Virgo. A number of imaging surveys (e.g., DESI and DECaLS) using the DECam wide field imager on the Blanco Telescope at CTIO have been photometrically calibrated and are now publicly available as part of the Legacy Surveys (Dey et al., 2019) and can be explored through NOIRLab’s Astro Data Lab. [^1] https://datalab.noirlab.edu/sia.php We retrieved data in the Legacy Survey DR10 release [^2] https://www.legacysurvey.org/dr10/description/ through the server at the National Energy Research Scientific Computing Center (NESRC). We combined the various archival images (“bricks”) for 23 target galaxies in the Jensen et al. (2021) sample and measured their $(g{-}z)$ colors. Uncertainties were determined by similarly combining the variance frames. The DECam Legacy Survey photometric uncertainties are 0.0073 mag for $g$ and $z$, which were combined in quadrature with each other and with the uncertainty in the foreground extinction correction from Schlafly & Finkbeiner (2011). The mean uncertainty in $(g{-}z)$ is 0.011 mag.
@@
Figure 1: Comparison of the PanSTARRS $(g{-}z)$ colors with those transformed from the DECam colors to PanSTARRS using Dey et al. (2019) for the seven calibrator galaxies with color data in both surveys. The black line denotes the one-to-one relation and is not a fit to the data.
@@
The eight galaxies for which JWST TRGB distances were individually measured (black points in the right half of Figure 2) provide a zero point good to 0.018 mag but a relatively weak constraint on the slope due to the limited range in color covered by the JWST sample (the slope is $1.7\pm 0.3$ in the PanSTARRS $(g{-}z)$ photometric system). To determine a more robust measurement of the slope, we adopted the mean cluster TRGB distance for the Virgo and Fornax clusters, and used those distances to compute $\overline{M}_{110}$ for a larger sample of HST Virgo and Fornax calibrators from Jensen et al. (2015) that include a wider range of galaxy color. When we include six additional galaxies and calculate $\overline{M}_{110}$ using mean TRGB cluster distances (see Table 1) we get a tighter value for the slope of $1.86\pm 0.16$ with a reduced $\chi_{\nu}^{2}=0.74$ per degree of freedom. The fact that the reduced $\chi_{\nu}^{2}$ is less than one suggests that the uncertainties in the absolute magnitude may be slightly overestimated, but there is still a 30% probability of measuring this value of $\chi_{\nu}^{2}$ for the number of degrees of freedom in our sample, indicating general consistency. Regardless of which TRGB distances (and slope) are used, the zero point at a color of $(g{-}z)=1.30$ is consistently $-2.76$ with an uncertainty below 0.02 mag.

### v2

@@ Abstract
The TRGB--SBF Project team is developing an independent distance ladder using a geometrical calibration of the tip of the red giant branch (TRGB) method in elliptical galaxies that can in turn be used to set the surface brightness fluctuation (SBF) distance scale independent of Cepheid variables and Type Ia supernovae (SNe Ia). The purpose of this project is to measure the local expansion rate of the universe independently of the methods that are most at odds with the theoretically-predicted value of the Hubble-Lemaître constant $H_{0}$, and therefore isolate the influence of potential systematic observational errors. In this paper, we use JWST TRGB distances calibrated using the megamaser galaxy NGC 4258 to determine a new Cepheid-independent SBF zero point for HST. This new calibration, along with improved optical color measurements from PanSTARRS and DECam, gives an updated value of $H_{0}=73.8\pm 0.7$ (statistical) ±$\pm 2.3$ (systematic) km s-1 Mpc-1 that is virtually identical to the SBF Hubble constant measured by Blakeslee et al. (2021).
@@ 1 Introduction
The current disagreement between the measured expansion rate of the universe and that inferred from the best fit version of the $\Lambda$CDM cosmological model derived from the cosmic microwave background fluctuations is now greater than 5-$\sigma$ (Planck Collaboration et al., 2020; Riess et al., 2022). Resolving this discrepancy (often referred to as the “Hubble tension”) is one of the most pressing issues in cosmology today. It is not yet clear if the explanation is an accumulation of various systematic errors in local distance measurements or if modifications to the $\Lambda$CDM model are required (e.g., Di Valentino et al., 2021, 2025).
Most measurements of $H_{0}$ in the relatively nearby universe favor values around 73 km s-1 Mpc-1 (see the recent compilation by Di Valentino et al. 2025). The SH0ES team has achieved the highest precision to date using four geometrical methods to anchor the Cepheid+SNe Ia distance ladder (Riess et al., 2022, 2024a). The Carnegie Chicago Hubble Program (CCHP) Team, in contrast, measured values of $H_{0}$ that are closer to the CMB+$\Lambda$CDM values (Freedman et al., 2024; Hoyt et al., 2025). The discrepancy between these two groups, using many of the same targets and similar methods, is most likely due to choices in the sample selection (Riess et al., 2024b) and not the geometrical anchors themselves. One of the primary purposes of this project is to test that conclusion by providing a measurement of $H_{0}$ that is completely independent of the Cepheid and SNe Ia techniques, and therefore to distinguish between systematic errors in the zero point calibrations and sample selection or analysis differences. If the Hubble tension between measurements from the CMB and $H_{0}$ measured locally persists with multiple independent techniques, then the possibility of additional physics beyond $\Lambda$CDM must be taken seriously.
@@
The TRGB--SBF Project seeks to reduce or even eliminate some sources of systematic uncertainty in the first rungs of the distance ladder by developing a new and independent distance scale based on a geometrical calibration of the old, metal-rich RGB population in massive elliptical galaxies. The new distance scale is completely independent of the Cepheid and SNe Ia distance measurements, and therefore avoids the systematic uncertainties associated with them. To do this, we are calibrating SBF distances measured in 14 nearby elliptical galaxies using TRGB distance measurements from the James Webb Space Telescope (Tully et al., 2023). The TRGB--SBF Project will eventually link geometrically-calibrated TRGB distances at 15–20 Mpc with SBF distances reaching out to perhaps 300 Mpc or more using the same telescope (JWST), camera (NIRCam), and filters to minimize systematic uncertainties for each rung of the ladder.
@@
SBF distances have already been measured in more than 370 galaxies out to about 100 Mpc using the Hubble Space Telescope (HST), including more than 220 in F110W using the WFC3/IR camera. In this paper, we use the TRGB distances to eight individual galaxies calibrated using the geometrical distance to the megamaser galaxy NGC 4258 (Reid et al., 2019; Anand et al., 2024a) to calibrate the HST SBF distances via TRGB. In the future, the Gaia satellite (Gaia Collaboration et al., 2017, 2018) will add additional geometrical anchors to the TRGB distance scale to further reduce the systematic calibration uncertainty. This study is the third paper in a series to establish an independent distance ladder based on Population II stars. The first two papers determined TRGB distances to multiple galaxies in the Fornax and Virgo clusters (Anand et al., 2024b, 2025). This paper presents the first step towards an SBF calibration that is independent of the Cepheid distance scale and connects JWST TRGB to the existing HST SBF distance measurements out to 100 Mpc.
@@ 2 Infrared Surface Brightness Fluctuations with HST
Almost all optical and infrared (IR) HST SBF measurements published in the last decade are based on the calibrations published by Blakeslee et al. (2009), Blakeslee et al. (2010), and Jensen et al. (2015, 2021). The foundation of SBF measurements used in these studies are the HST Cepheid distances to seven galaxies with ground-based SBF distances measured by Tonry et al. (2001) and Blakeslee et al. (2002), combined with a large number of overlapping ground-based and HST SBF measurements in the Virgo and Fornax galaxy clusters (see Blakeslee et al., 2010, for direct comparisons). These large surveys provided the statistical foundation for SBF and showed that the intrinsic scatter in the absolute fluctuation magnitudes (and therefore the SBF distance moduli) is  ${\sim\,}0.06$ mag (optical) to  ${\sim\,}0.1$ mag (near-IR), after correction for stellar population variations using optical colors (Blakeslee et al., 2009; Jensen et al., 2015). The current project seeks to replace the Cepheid calibration with TRGB for the purpose of isolating systematic uncertainties in the various distance measurement techniques.
@@
Fluctuations are more prominent at near-IR wavelengths than in the optical, and are measurable to ${\gtrsim}80$ Mpc in a single HST orbit. The SBF method is much more efficient than Cepheids or SNe Ia for measuring distances out to 100 Mpc because only one imaging observation is required, and SBF achieves comparable precision in much less observing time. The majority of the SBF distances beyond 40 Mpc useful for measuring $H_{0}$ were made using NICMOS (Jensen et al., 2001) and WFC3/IR (Cantiello et al., 2018; Blakeslee et al., 2021). The most recent WFC3/IR F110W SBF calibration was determined using 11 galaxies in the Fornax and Virgo galaxy clusters, and was corrected for stellar population variations using ACS $(g_{475}{-}z_{850})$ colors (Jensen et al., 2015). Updates to the Jensen et al. (2015) calibration to include better Galactic foreground extinction corrections and an improved distance to the LMC were made as described by Cantiello et al. (2018) and Jensen et al. (2021).
@@
The majority of the HST IR SBF observations published between 2015 and 2021 were part of HST programs GO-14219 (PI J. Blakeslee), GO-14654 (PI P. Milne), and GO-15625 (PI J. Blakeslee). Jensen et al. (2021) published SBF distances to 63 galaxies from these programs, including 25 SNe Ia host galaxies. Three more HST programs have recently been completed. GO-16262 (PI R. Tully), GO-17436 (PI J. Jensen), and GO-17446 (PI P. Milne) have already observed ${\gtrsim}164$ more galaxies out to ${\gtrsim}85$ Mpc, a sample that more than doubles the number of SNe Ia host galaxies with high-quality SBF distances (now 51). To make the most of these new datasets, in this paper we update the HST SBF calibration using the new JWST TRGB distances.
@@
The SBF distances published by Jensen et al. (2021) were used by Blakeslee et al. (2021) to measure a value of the Hubble constant of $H_{0}=73.3\pm 0.7$ (statistical) ±$\pm 2.4$ (systematic)  km s-1 Mpc-1. The subsample of 25 early-type SNe Ia host galaxies with HST SBF distances provided an alternative calibration of SNe Ia in the Hubble flow: Garnavich et al. (2023) measured a value of $H_{0}=74.6\pm 0.9$ (statistical) ±$\pm 2.7$ (systematic) km s-1 Mpc-1 using SBF distances as an intermediate step between Cepheids and SNe Ia. Garnavich et al. (2023) used the Pantheon+ (Scolnic et al., 2022) light curves to make the SNe Ia distance measurements. When the light curve fitting parameters were corrected to better match the fast-declining SNe Ia preferentially found in the early-type galaxies, the value of $H_{0}$ became $73.3\pm 1.0\pm 2.7$ km s-1 Mpc-1.
@@ 3 Re-calibrating the HST WFC3/IR SBF Distance Scale using TRGB
JWST observations of the TRGB in 14 nearby elliptical galaxies (GO-3055; Tully et al. 2023) provide a new and direct calibration of IR SBF based on the geometrical megamaser distance to NGC 4258 (Anand et al., 2024a; Reid et al., 2019). We will eventually use JWST to determine $H_{0}$ directly by measuring SBF distances out to  ${\sim}\,300$ Mpc, but building a database of distant SBF measurements will take time, and may never reach the quantities already observed with HST over the last 30 years (more than $370$ galaxies, including SBF distances from NICMOS, ACS, and WFC3/IR). To leverage the full potential of the HST database, we need to connect the new JWST TRGB distances to the existing HST SBF dataset. Three of the JWST TRGB target galaxies in the Fornax cluster and five in the Virgo cluster have also been observed with WFC3/IR, as shown in Table 1. The new JWST TRGB measurements of these galaxies (Anand et al., 2024b, 2025) provide us with an opportunity to reduce systematic uncertainties in SBF distances originating from the Cepheid calibration. Using a direct geometrical calibration of TRGB from NGC 4258 (Reid et al., 2019; Anand et al., 2024a) and SBF measurements in the same galaxies, using the same telescope, camera, and filters, we can eliminate some sources of uncertainty (e.g., Cepheid metallicity corrections in different galaxy types and positional offsets between spirals and ellipticals in the calibration clusters).
@@ 3.1 Updated Calibration
In this paper we re-calibrated the HST IR SBF distance scale by replacing the Cepheid-based optical SBF distances from Jensen et al. (2015) with the new JWST-based TRGB distances to the eight galaxies in common between the HST and JWST SBF calibration data sets. The SBF calibration of the absolute fluctuation magnitude in the F110W filter, denoted with the symbol $\overline{M}_{110}$, depends on the distance scale zero point and a slope with galaxy color, which corrects $\overline{M}_{110}$ for variations in stellar population age and metallicity (e.g., Jensen et al., 2003). Jensen et al. (2015) found that the rms scatter among the calibration galaxies was 25% smaller when using the average distance to the cluster instead of individual optical SBF distances. [^1] Whether or not the individual optical SBF distances will give a better calibration depends on the magnitude of the uncertainties relative to the cluster depth. When individual observational uncertainties are larger than the cluster depth, averaging first produces a lower rms in the calibration scatter; when they are smaller than the cluster depth, then using the individual distances will result in a lower scatter. In the case of the ACS SBF measurements in Virgo and Fornax, the individual uncertainties and the cluster depth were comparable (see Jensen et al. 2015 for details). We can do the same here, and by using the mean cluster distances from the JWST TRGB measurements we include four more galaxies in the Fornax cluster and two more in the Virgo cluster, for a total of 14 galaxies in our new calibration (see Table 1).
@@
footnotetext: (4) SBF magnitudes and optical colors have been corrected for Galactic extinction using Schlafly & Finkbeiner (2011) and the NED extinction calculator https://doi.org/10.26132/NED5 (catalog DOI: 10.26132/NED5). Extinction corrected values are indicated with a subscript 0. All fluctuation magnitudes and colors were measured in two annular region around each galaxy from 8.2 to 16.4 arcsec and 16.4 to 32.8 arcsec in radius. The values shown are the weighted averages of the two measurements.
@@
The direct overlap between JWST (TRGB) and HST (Cepheid) calibration samples includes three galaxies in the Fornax cluster and five in the Virgo cluster (including NGC 4636 and NGC 4697 in peripheral in-falling regions). Using the combined average TRGB distance modulus to the Virgo cluster of $31.043\pm 0.034$ (statsitical) ±$\pm 0.063$ (systematic) mag (Anand et al., 2024b, 2025) and the Cepheid-based ACSVCS survey average distance modulus of $31.092\pm 0.013$ mag (Mei et al., 2007; Blakeslee et al., 2009), which was used to calibrate the previous IR SBF measurements (Jensen et al., 2021), the weighted mean difference is $\Delta\mu_{\rm TRGB-Ceph}=-0.067\pm 0.031$ mag (Anand et al., 2025). For Fornax, the difference between the mean ACS SBF distance modulus of 31.51 mag (Blakeslee et al., 2009) used for the Jensen et al. (2015) IR SBF calibration and the mean distance from three JWST TRGB measurements gives a difference of $\Delta\mu=-0.086\pm 0.077$ mag, where the uncertainty is the result of the scatter among the three galaxies.
@@
The SBF measurements were updated since the original calibration published by Jensen et al. (2015), and for this study we have included three additional calibration galaxies (NGC 4552, NGC 4636, and NGC 4649). The new apparent SBF magnitude $\overline{m}_{110}$ measurements for the calibrators are consistent with the procedures we used to measure $\overline{m}_{110}$ in the sample from which we derived $H_{0}$ (Jensen et al., 2021; Blakeslee et al., 2021). The average difference between the SBF magnitudes $\overline{m}_{110}$ from the Jensen et al. (2015) calibration paper and the present values for the five calibrator galaxies in common is 0.072 mag with a standard deviation of 0.03 mag. Jensen et al. (2021) introduced a number of improvements to the SBF analysis (e.g., improvements to the PSF normalization, modernization of the LMC zero point calibration for Cepheids, and improvements to the procedure to remove SBF signal resulting from undetected globular clusters) that were previously applied to the sample used to determine $H_{0}$ (Blakeslee et al., 2021), and those have been applied to these new measurements. We previously identified an offset of $0.05\pm 0.02$ mag to the 2015 calibration values due to the PSF normalization (Cantiello et al., 2018). Given the small sample size in this study, the 0.072 mag offset is considered to be statistically consistent with the previously measured 0.05 mag offset. Further assessment of the reliability of the $\overline{m}_{110}$ measurements will require additional overlap between calibration samples.
@@
Optical galaxy colors are critical for the SBF method as they are necessary to correct SBF amplitudes for intrinsic variations in the ages and metallicities of the stellar populations in elliptical galaxies (Jensen et al., 2003; Blakeslee et al., 2009; Jensen et al., 2015). The original IR SBF calibration was based on the ACS $(g_{475}{-}z_{850})$ colors, which are not generally available for the vast majority of the WFC3/IR observations in the HST archive. Jensen et al. (2021) transformed PanSTARRS (PS) $(g{-}z)$ colors (Chambers et al., 2019) to the ACS photometric frame for computing distances. For this study, we remeasured all the PanSTARRS $(g{-}z)$ colors and determined the SBF calibration directly in that system. We found that for galaxies with significant color gradients, the PanSTARRS $(g{-}z)$ colors used by Jensen et al. (2021) for computing SBF distances—measured in the high-$S/N$ ratio central regions—did not always match the colors in the regions where the SBF was measured. For typical massive galaxies with slightly redder cores and bluer halos, correcting the colors to match the SBF regions leads to larger distances, by as much as 5% in some cases, and hence this revision tends to reduce the value of $H_{0}$. For the 61 galaxies in the Jensen et al. (2021) dataset with PanSTARRS photometry, which are used again in this paper to recompute $H_{0}$, the color shift was measured to be $-0.023\pm 0.019$ mag (bluer). The average photometric uncertainty in $(g{-}z)$ is 0.02 mag for PanSTARRS, which includes the uncertainty in the extinction correction (Schlafly & Finkbeiner, 2011) added in quadrature. Given that the historical slope of the relationship between SBF $\overline{M}_{110}$ and $(g{-}z)$ is  ${\sim\,}2$ (Jensen et al., 2015), the expected shift in distance modulus is approximately 0.05 mag or 2.3% in distance, counteracting in part the increase in $H_{0}$ expected from the new TRGB zero point.
@@
The original SBF calibration of Jensen et al. (2015) included equal numbers of Fornax and Virgo cluster galaxies, which were calibrated using ACS optical photometry. The JWST TRGB sample includes three galaxies in the Fornax cluster. The PanSTARRS-1 survey does not reach far enough south to include Fornax, so the current TRGB calibration is limited to five galaxies in the Virgo cluster. Surveys using DECam on the Blanco Telescope at CTIO provide a high-$S/N$ alternative to PanSTARRS for both the Fornax and Virgo clusters. A number of imaging surveys (e.g., DESI and DECaLS) using DECam are now publicly available as part of the Legacy Surveys (Dey et al., 2019). [^2] https://datalab.noirlab.edu/sia.php We retrieved data in the Legacy Survey DR10 release [^3] https://www.legacysurvey.org/dr10/description/ through the server at the National Energy Research Scientific Computing Center (NESRC). We combined the archival images (“bricks”) for 23 target galaxies in the Jensen et al. (2021) sample and measured their $(g{-}z)$ colors. Uncertainties were determined by similarly combining variance frames. The DECam Legacy Survey photometric uncertainties are 0.0073 mag for $g$ and $z$, which were combined in quadrature with each other and with the uncertainty in the foreground extinction correction from Schlafly & Finkbeiner (2011). The mean uncertainty in $(g{-}z)$ is 0.011 mag.
@@
Figure 1: Comparison of the extinction-corrected PanSTARRS $(g{-}z)$0 colors with those transformed from the DECam colors to PanSTARRS using Dey et al. (2019) for the seven calibrator galaxies with color data in both surveys. The black line denotes the one-to-one relation and is not a fit to the data.
@@
The eight galaxies for which JWST TRGB distances were individually measured (black points in the right half of Figure 2) provide a zero point with an accuracy of 0.018 mag but a relatively weak constraint on the slope due to the limited range in color covered by the JWST sample (the slope is $1.7\pm 0.3$ in the PanSTARRS $(g{-}z)$ photometric system). To determine a more robust measurement of the slope, we adopted the mean cluster TRGB distance for the Virgo and Fornax clusters, and used those distances to compute $\overline{M}_{110}$ for a larger sample of HST Virgo and Fornax calibrators from Jensen et al. (2015) that include a wider range of galaxy color. When we include six additional galaxies and calculate $\overline{M}_{110}$ using mean TRGB cluster distances (see Table 1) we get a tighter value for the slope of $1.86\pm 0.16$ with a reduced $\chi_{\nu}^{2}=0.74$ per degree of freedom. The fact that the reduced $\chi_{\nu}^{2}$ is less than one suggests that the uncertainties in the absolute magnitude may be slightly overestimated, but there is still a 30% probability of measuring this value of $\chi_{\nu}^{2}$ for the number of degrees of freedom in our sample, indicating general consistency. Regardless of which TRGB distances (and slope) are used, the zero point at a color of $(g{-}z)=1.30$ is consistently $-2.76$ with an uncertainty below 0.02 mag.